Find the volume of the sphere in terms of pi in inches cubed. 15 inches is the diameter. (Type the number and then type pi. Do not round your answer. )
Answer by Boreal(15235)

V=(4/3)*pi*r^3
=(4/3)(7.5)^3*pi
=562.5*pi in^3 volume